OpenStudy (anonymous):

Find \(p,q\) to the expression \(A = p (cos^{8}x-sin^{8}x) + 4(cos^{6}x-2sin^{6}x) + qsin^{4}x\) does not depend upon \(x\)

OpenStudy (dan815):

okay if u dont want it to depend on x, we have to compress this into sin^2 x + cos^2 (x) identity or something similar

OpenStudy (dan815):

... is this what all korean problems in school are like

OpenStudy (dan815):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html here u go have fun

ganeshie8 (ganeshie8):

may be try messing wid exponents + binomial thm instead : \(\cos x = \dfrac{e^{-ix} + e^{ix}}{2}\) \(\sin x = \dfrac{e^{-ix} - e^{ix}}{2}\)

OpenStudy (dan815):

complex numbers are really amazing

OpenStudy (dan815):

i shud play with them more often

ganeshie8 (ganeshie8):

dan lookup star polygons using complex numbers... i checked that few days back because of @Kainui and it was really wonderful !

OpenStudy (dan815):

do u mean the fractcal pattern generation with complex numbers?

ganeshie8 (ganeshie8):

Yes more specifically about iterations

OpenStudy (dan815):

ok cool

OpenStudy (kainui):

Interesting that you mention this @ganeshie8 because I was just thinking about that and looking at star polygons today and thinking about applying them to making knots. Just a little playing around and I discovered that I could make some pretty cool looking knots and use a little bit of complex numbers to play around with them.